# Fast Trac: A Parameter-Free Optimizer for Lifelong Reinforcement Learning

### v2

@@ Appendix C Gravity Based Distribution Shifts
Appendix D Scaling-Value Convergence
@@
The convergence of the scaling parameter $S_{t+1}$ observed across the Procgen and Gym Control environments, as depicted in Figures 13 and 15, reflects a good scaling value that effectively determines the strength of regularization towards the initialization points, yielding robust empirical outcomes in lifelong RL settings. Interestingly, in Procgen environments, this converged scaling value exhibits consistency across various games, typically hovering between 0.02 and 0.03, as shown in Figure 13. In contrast, in the Gym Control environments, the scaling values are lower, ranging between 0.005 and 0.01, as illustrated in Figure 15.
Appendix E Comparison to Mechanic
In our analysis, we extend the examination to other OCO-based optimizers within the lifelong RL framework. Figure 16 presents a comparative assessment of Trac PPO and Mechanic PPO (Cutkosky et al., 2023) for Procgen and Gym Control tasks. The results indicate that while Mechanic effectively mitigates plasticity loss and adapts quickly to new distribution shifts, it slightly underperforms in comparison to Trac. This suggests potential for further explorations and a broader evaluation of OCO-based optimizers in lifelong learning environments.
Figure 16: Reward in the Procgen and Control environments with both Trac PPO and Mechanic PPO. Mechanic PPO performs similarly to Trac PPO, although slightly underperforming Trac PPO.
Appendix F Experimental Setup
@@
The Procgen and Atari experiments were conducted with 8 seeds/runs, while the Gym Control experiments utilized 25 seeds/runs. The exception was in the $L_{2}$ initialization experiments, which used 15 seeds/runs per regularization strength. In Figures 4, 5, 6, 7, 9, 10, 11, 12, the plotted lines represent the mean of all of the mean episode rewards from the different seeds/runs, and the shaded error bands indicate the standard deviation of all of the mean episode rewards from the different seeds/runs.

### v3

@@ Appendix D Gravity Based Distribution Shifts
Appendix E LayerNorm, Plasticity Injection, and Weight Decay
To evaluate Trac alongside other methods that aim to mitigate plasticity loss, we compare it against LayerNorm (Lyle et al., 2023), Plasticity Injection (Nikishin et al., 2024), and tuning weight decay (Lyle et al., 2024).
As discussed in Section 5, we confirm that both layer normalization and plasticity injection (applied at the start of every distribution shift) (Nikishin et al., 2024; Lyle et al., 2023) are effective in reducing plasticity loss (Figure 15). While these methods help slow the decline in performance due to plasticity loss, Trac consistently outperforms them across the three Gym Control environments. Importantly, because Trac is an optimizer, it can be combined with layer normalization, and doing so resulted in the best performance gains in our Control setups.
Tuning weight decay: In addition to LayerNorm and Plasticity Injection, we also evaluated the effects of tuning weight decay using PyTorch’s AdamW optimizer. We conducted a hyperparameter sweep across three control environments with 15 seeds for each of the following weight decay values: 0.0001, 0.001, 0.01, 0.1, 1.0, 5.0, 10.0, 15.0, and 50.0. Figure 16 presents the average normalized reward for each weight decay value over 15 seeds and 3000 timesteps, compared to Trac.
The results indicate that while tuning weight decay with Adam does provide some benefit, these values consistently underperform in comparison to Trac across all three control environments. Figure 16 plots the performance of the best-performing weight decay value with Adam over 10 distribution shifts in the control environments. We observe that weight decay values are highly sensitive to the specific environment and the nature of the distribution shift.
Interestingly, in our initial experiments, we set the weight decay to zero, yet Trac still outperformed Adam with various weight decay values. This suggests that while weight decay can mitigate plasticity loss to some extent, it does not match the overall effectiveness of Trac.
Figure 15: Performance comparison of plasticity loss mitigation techniques across Gym Control environments. Both layer normalization and plasticity injection reduce plasticity loss when applied with Adam. Trac outperforms both layer norm Adam and plasticity injection Adam, with the combination of layer norm and Trac achieving the highest performance.
Figure 16: Effect of weight decay on performance in the three Gym Control environments. Bar plots show the average normalized rewards over 25 seeds for different weight decay values using Adam across 3000 timesteps, compared to Trac with no weight decay.
Appendix F Scaling-Value Convergence
@@
The convergence of the scaling parameter $S_{t+1}$ observed across the Procgen and Gym Control environments, as depicted in Figures 17 and 19, reflects a good scaling value that effectively determines the strength of regularization towards the initialization points, yielding robust empirical outcomes in lifelong RL settings. Interestingly, in Procgen environments, this converged scaling value exhibits consistency across various games, typically hovering between 0.02 and 0.03, as shown in Figure 17. In contrast, in the Gym Control environments, the scaling values are lower, ranging between 0.005 and 0.01, as illustrated in Figure 19.
Appendix G Comparison to Mechanic
In our analysis, we extend the examination to other OCO-based optimizers within the lifelong RL setup. Table 3 presents a comparative assessment of Trac PPO and Mechanic PPO (Cutkosky et al., 2023) for the lifelong Gym Control tasks (with 300 seed runs). The p-values were calculated using two-sample t-tests to test the hypothesis that the means between Trac and Mechanic are the same (Null Hypothesis, $H_{0}$) against the alternative hypothesis that they are different (Alternative Hypothesis, $H_{1}$). The results indicate that while Mechanic effectively mitigates plasticity loss and adapts quickly to new distribution shifts, it slightly underperforms in comparison to Trac.
Table 3: Performance comparison between Trac and Mechanic across three Gym Control environments. The mean, standard error, and p-values reflect the performance over multiple runs, with bolded values highlighting Trac’s superior results.
Task Method Mean Std Error p-value
LunarLander-v2 Trac 0.6018 0.0036 0.0000
Mechanic 0.5755 0.0027
CartPole-v1 Trac 0.3518 0.0244 0.0021
Mechanic 0.3008 0.0230
Acrobot-v1 Trac 0.7044 0.0221 0.0000
Mechanic 0.6396 0.0239
Appendix H Experimental Setup
@@
The Procgen and Atari experiments were conducted with 8 seeds/runs, while the Gym Control experiments utilized 25 seeds/runs (with the exception of the Mechanic experiments in Table 3 which utilized 300 seeds). The exception was in the $L_{2}$ initialization experiments, which used 15 seeds/runs per regularization strength. In Figures 4, 5, 6, 7, 9, 10, 11, 15, 12, 14, the plotted lines represent the mean of all of the mean episode rewards from the different seeds/runs, and the shaded error bands indicate the standard deviation of all of the mean episode rewards from the different seeds/runs.